OpenStudy (anonymous):

5(1-t)<4(3-t). Trying to find if the 5(1-t) is less than or equal to 4(3-t)

OpenStudy (anonymous):

\[5-5t \le 12-4t\]\[-5t \le 7-4t\]\[-t \le 7\]\[t \ge -7\]

OpenStudy (anonymous):

Thank you